If x and y are arbitrary real numbers with x < y, prove that there exists at least one irrational number z satisfying x < z < y, and hence infinitely many.


1
Expert's answer
2021-01-25T03:29:23-0500

Here, we have "x',y'\in\R," such that "x'<y'."


Now, we already know that there exists a rational number "q" such that "x'<q<y'."


Now, let "n\in\R," so "\sqrt{n}\in\R" .


Dividing all sides by "\sqrt{2}" :-



"\frac{x'}{\sqrt{2}}<\frac{q}{\sqrt{2}}<\frac{y'}{\sqrt{2}}"

Now, as "q\in\mathbb{Q}," so "\frac{q}{\sqrt{2}}=z" "\in\mathbb{Q^c}" .


And "x',y'\in\R", so we have "\frac{x'}{\sqrt{2}},\frac{y'}{\sqrt{2}}" "\in\R"


So,

"x<z<y"

where, "x=\frac{x'}{\sqrt{2}}" and so on.


So, we have an irrational number between "x,y" such that "x<z<y" .


Now, we took "\sqrt{2}" as an example, but we can replace by any "n\in\R" . Note that some "n=4,16" are rational, but we have infinitely many "n\in\R" (as "\R" is an infinite set), such that "\frac{q}{\sqrt{2}}" "\in\mathbb{Q^c}" .